solve the system {
x-y-2z=1
x-5y+2z=5
2x-3y-4z=2

Answer by amarjeeth123(569)

x-y-2z=1.................equation 1
x-5y+2z=5................equation 2
2x-3y-4z=2...............equation 3
Adding equations 1 and 2 we get,
2x-6y=6..................equation 4
Multiplying equation 2 by 2 we get,
2x-10y+4z=10.............equation 5
Adding equations 3 and 5 we get,
4x-13y=12................equation 6
Multiplying equation 4 by 2 we get,
4x-12y=12................equation 7
Subtracting equation 7 from equation 6 we get,
y=0
x=3
z=1
The solution is x=3,y=0 and z=1.